# On the Fairness of Privacy Protection: Measuring and Mitigating the Disparity of Group Privacy Risks for Differentially Private Machine Learning

### v1

@@ Abstract
While significant progress has been made in conventional fairness-aware machine learning (ML) and differentially private ML (DPML), the fairness of privacy protection across groups remains underexplored. Existing studies have proposed methods to assess group privacy risks, but these are based on the average-case privacy risks of data records. Such approaches may underestimate the group privacy risks, thereby potentially underestimating the disparity across group privacy risks. Moreover, the current method for assessing the worst-case privacy risks of data records is time-consuming, limiting their practical applicability. To address these limitations, we introduce a novel membership inference game that can efficiently audit the approximate worst-case privacy risks of data records. Experimental results demonstrate that our method provides a more stringent measurement of group privacy risks, yielding a reliable assessment of the disparity in group privacy risks. Furthermore, to promote privacy protection fairness in DPML, we enhance the standard DP-SGD algorithm with an adaptive group-specific gradient clipping strategy, inspired by the design of canaries in differential privacy auditing studies. Extensive experiments confirm that our algorithm effectively reduces the disparity in group privacy risks, thereby enhancing the fairness of privacy protection in DPML.
@@ 1 Introduction
Artificial intelligence (AI), particularly machine learning (ML), has been widely adopted across various sectors, augmenting and even replacing human decision-making. However, its growing integration in critical domains like healthcare, finance, and judiciary has raised critical concerns, including data privacy breaches, algorithmic biases, lack of explainability, security vulnerabilities etc. [10]. Among these ethical issues and risks, privacy and fairness have emerged as two pivotal and widely discussed challenges [8], attracting substantial attention from the research community.
@@
Research in privacy protection and fairness has made significant strides independently, and the intersection of these two critical issues has also gained considerable attention. Some studies aim to achieve both privacy protection and outcome fairness simultaneously in ML models [26, 11, 5, 25, 15]. Other works explore how privacy mechanisms affect the outcome fairness [3] and propose methods to mitigate the unfairness introduced by such mechanisms [27, 24, 9]. However, whether AI systems can provide equal privacy protections to different groups is also a noteworthy yet understudied problem at the intersection of fairness and privacy. As highlighted in [8], this raises an essential yet still underexplored and insufficiently addressed question: Do AI systems offer fair or equitable privacy protections across groups? This question raises both ethical and practical concerns, as certain groups may face disproportionately higher privacy leakage risks, violating principles of fairness and equality.
@@
To address this issue, it is essential first to provide a rigorous answer to the question. Prior studies have empirically examined whether privacy leakage risks are evenly distributed across groups, and their auditing methods for group privacy risk rely on averaging the performance across data points within each group under membership inference attacks (MIAs) [4, 28]. The MIAs employed in these studies are formulated based on the membership inference game (MIG) introduced in [30], which captures the average behavior across all data points. However, such an average-case MIG may obscure the heterogeneity of individual privacy risks within groups and potentially underestimate the privacy leakage risks faced by certain groups. This, in turn, may lead to inaccurate and unreliable measurements of inter-group disparities in privacy risk. Consequently, the issue of privacy inequality may not be adequately uncovered.
@@
Measuring. Therefore, we aim to provide a tighter measurement of the privacy risk of data points and then analyze the disparities in privacy risk across different groups. To this end, we can leverage the leave-one-out attack (LOOA), which is formulated based on the worst-case MIG proposed by [29] and is capable of estimating the worst-case privacy leakage risk for each data point. However, the computational cost of LOOA is prohibitively high, rendering its practical implementation nearly infeasible. To address this challenge, we propose an approximate version of the worst-case MIG to efficiently audit the approximate worst-case privacy risk of individual data points. Our experiments demonstrate that: 1) The attack simulating our proposed MIG can achieve comparable performance to LOOA as the number of attack rounds increases; 2) The individual privacy risks evaluated by our method can be reliable for assessing group privacy risk.
@@
Then we define a fairness metric to quantify the degree of privacy unfairness (i.e., the disparity of privacy risk across groups). Through experiments, we clearly demonstrate that our auditing method significantly outperforms previous auditing methods under the same conditions. Specifically, our method reveals greater group privacy risk and more effectively captures privacy inequality. Consistent with prior research [4, 28], we find that existing ML algorithms exhibit significant unfairness in privacy risks across groups. While differentially private ML (DPML) algorithms can bind the magnitude of privacy risk disparities between groups, a certain degree of disparity still persists.
@@
Mitigating. Upon providing a more thorough answer to the above question, we seek to alleviate this issue. Inspired by the design of canaries in DP auditing studies [18, 2, 23], we confirm that groups with larger gradient norms under training process—indicating greater contributions to model updates—are more prone to higher privacy leakage risks. Building on this insight, we enhance the existing DPML algorithm by adaptively setting group-specific gradient clipping norms. Extensive experimental results demonstrate that our algorithm effectively mitigates the disparity of group privacy risk, promoting the ethical and effective deployment of AI systems.
@@ 2.1 Differential privacy
Differential Privacy (DP), proposed by [7], is a privacy framework designed to address privacy leakage. It has become the predominant method for ensuring algorithmic privacy [20]. In the following, we introduce the approximate ($\epsilon,\delta$)-DP definition.
Definition 1 ($(\epsilon,\delta)$-Differential Privacy [6]).
@@
DP-SGD [1] is a widely adopted algorithm in DPML [20]. It integrates DP concepts with stochastic gradient descent (SGD). This integration ensures model privacy by employing gradient clipping and noise addition within the SGD framework, adhering to the $(\epsilon,\delta)$-DP definition. The pseudocode of DP-SGD is shown in Algo. 2.
@@ 2.2 Black-box member inference attacks
The goal of membership inference attacks (MIAs) is to determine whether a specific data record is part of the training dataset. We focus on a black-box setting, where the adversary only has access to model outputs, reflecting a more realistic scenario where the training process is inaccessible [2]. In black-box MIAs, the adversary infers membership by analyzing the model’s output behavior, typically using the sample’s output loss as an inference score or decision basis, relying on the observation that models tend to show smaller losses for training samples [30].
@@
MIGs conceptualize MIAs as inference games between a privacy auditor (i.e., the adversary) and a challenger. MIAs are typically carried out by simulating the MIGs through multiple rounds of random experiment. Various definitions of MIGs have been proposed, each designed to capture different aspects of privacy risk [29].
Most MIAs follow the average-case MIG framework [30] (see Def.6 in App.A)), which evaluates the vulnerability of a target model to the adversary, emphasizing the average behavior across data points [29]. A common strategy formalized under this framework is the global attack (GA), where a single inference threshold is determined based on the aggregate behavior of all data points in a given round [30, 21]. The group-based attack (GBA) extends GA by assigning a distinct threshold to each group, enabling a more fine-grained analysis of group-level privacy risks and offering deeper insight into disparities across demographic partitions [4, 28].
Nonetheless, such average-case MIAs fail to capture worst-case privacy risks for individual data records. The worst-case MIG  [29] (see Def.7 in App.A) addresses this limitation by evaluating the maximum risk a single record may encounter. A concrete instance is the Leave-One-Out Attack (LOOA), which independently evaluates each data point’s worst-case exposure, aligning closely with the principles of DP. However, LOOA is computationally intensive, as it requires evaluating each record separately.
@@
Privacy auditing is designed to proactively assess privacy risks and quantify potential leakage, typically during the model development phase. In contrast, MIAs are conducted post-deployment by adversaries aiming to exploit trained models. Privacy auditing uses MIAs for evaluation but with more background knowledge for the adversary, including access to the original dataset and knowledge of the optimal threshold. This setup simulates worst-case scenarios, enabling rigorous assessment of privacy leakage risks [4]. In this work, we focus on privacy auditing to systematically evaluate privacy vulnerabilities. Privacy auditing using the GA (PA-GA)[30], the GBA (PA-GBA) [4, 28], and the LOOA (PA-LOOA) [18, 2] for evaluating privacy leakage risks is detailed in Algos. 34, and 5 of App. A, respectively.
@@ 2.3 The fairness of privacy
Few studies have explored fairness in the context of privacy, and those that do vary in their research focus. For instance, one study finds that fairness-aware algorithms can exacerbate disparities in privacy leakage across groups [4]. Meanwhile, other research highlights that ML algorithms exhibit significant group-level disparities in privacy leakage, and that DPML algorithms can help reduce these disparities [28]. These auditing mechanisms typically rely on average-case attacks, which assess privacy risks based on the average behavior of data points. While effective at capturing general trends, such approaches may overlook the nuanced privacy risks faced by individual samples, potentially concealing disparities between groups.
In our work, we aim to more precisely evaluate group-level privacy risks compared to prior studies. To ensure fair comparison, we adopt the same attack metrics as in [4, 28], including inference scores based on per-example loss and attack success rates.
@@ 3.1 Approximate worst-case privacy auditing
We validate our hypothesis through practical experiments using the widely used MNIST dataset, training a Convolutional Neural Network (CNN) with SGD. Due to computational limitations, we randomly select 60 samples per class, totaling 600 samples to audit for PA-LOOA and PA-ALOOA. We evaluate the attacker’s performance using the accuracy metric (i.e., attack success rate) from [22], which measures the agreement between the adversary’s guesses and the actual status. Instead of evaluating overall accuracy, we compute individual accuracy for each data point. Detailed experimental settings and additional results are provided in the App. B.
@@ 3.2 Definition of group privacy risk parity
We assess the privacy leakage risk of data points using the attacker’s membership advantage, following prior work [28, 30]. Below, we formally define the notion of individual privacy risk (IPR) as applied to a single data record in this study.
@@
This formulation quantifies the adversary’s normalized advantage over random guessing. Building on the concept of IPR, we can extend it to define Group Privacy Risk (GPR) as in  [4].
@@ 3.3 Comparison with privacy auditing by average-case attacks
We compare the GPR and the GPRP metrics measured by our auditing method, PA-ALOOA, with those obtained from privacy auditing by average-case attacks (PA-ACAs) in previous studies: PA-GA [28] and PA-GBA [4]. To ensure a fair comparison, all three methods are configured identically, keeping the training dataset and model consistent across each repeated experiment. Specifically, for PA-ALOOA, we set the number of audit samples to $m=n$, following the same setting used in PA-ACAs. This setup ensures consistency and reflects a real-world scenario in which the privacy risk of every training sample is audited. The key distinction lies in threshold determination: PA-ACAs computes thresholds based on the aggregate behavior of multiple data points, whereas PA-ALOOA assigns a unique threshold to each sample, based on its behavior across all repeated experiments.
@@
Consistent with prior studies on privacy and fairness [3, 27, 9], our foundational analysis focuses on the MNIST dataset. Detailed experimental configurations and supplementary results for other datasets are included in the App. C. The results of the GPR and GPRP metrics on the MNIST dataset are presented in Fig. 3 and Tab. 1, with all values reported in percentage points for clarity. As shown in Fig. 3, the three auditing methods exhibit consistent patterns in the distribution of privacy risks across groups. Among them, PA-ALOOA consistently yields significantly higher GPR values across all model architectures compared to the other two auditing methods. Specifically, Tab. 1 shows that for the CNN model, the GPRP value obtained by PA-GA and PA-GBA suggests that DP-SGD results in higher disparity than standard SGD. This observation contradicts the conclusion in [28], which asserts that DPML algorithms should be able to bind the disparity of GPR relative to non-private counterparts. Such inconsistency implies that PA-ACAs underestimate the GPR, leading to inaccurate measurements and incorrect conclusions. In summary, PA-ALOOA offers a more rigorous means of capturing privacy risk, and thus providing more reliable evaluations of privacy unfairness across groups.
@@ 4.1 Experimental observations
In DP auditing literature, canaries are often created as mislabeled samples [18, 2, 23]. These samples generate larger gradient values during model training, which contribute more to parameter updates, thereby increasing the likelihood of being memorized by the model. Motivated by the design of canaries, we hypothesize that during training, the larger a group’s contribution to the gradient, the more likely the model is to memorize that group. Thus, groups with larger contributions are expected to face a higher privacy leakage risk compared to those with smaller contributions.
@@ 4.2 Design of mitigation algorithm for DP-SGD
As shown in Algo. 1, our proposed algorithm, DP-SGD-Scale (abbreviated as DP-SGD-S), differs from the standard DP-SGD in Lines 6–14. In each iteration, it estimates the relative contribution of each group’s samples to the overall gradient and uses this information to adaptively adjust the clipping bound for each group. To preserve privacy, we add noise to the clipped group-level gradient statistics used for computing the group-specific clipping bounds $C^{k}$. This additional privacy cost is incorporated into the overall privacy accounting via the composition theorem [1, 27]. Although different groups are assigned distinct clipping bounds, we conservatively bound the sensitivity of the final aggregated gradient by $\max_{k}C^{k}$, ensuring that the overall mechanism satisfies $(\epsilon,\delta)$-DP in the same sense as DP-SGD. Following prior work [27, 9], we set $\sigma_{1}\approx 10\sigma_{2}$ so that the privacy cost of computing $C^{k}$ is negligible relative to the total privacy budget.

### v3

@@ Abstract
While significant progress has been made in conventional fairness-aware machine learning (ML) and differentially private ML (DPML), the fairness of privacy protection across groups remains underexplored. Existing studies have proposed methods to assess group privacy risks, but these are based on the average-case privacy risks of data records. Such approaches may underestimate the group privacy risks, thereby potentially underestimating the disparity across group privacy risks. Moreover, the current method for assessing the worst-case privacy risks of data records is time-consuming, limiting their practical applicability. To address these limitations, we introduce a novel membership inference game that can efficiently audit the approximate worst-case privacy risks of data records. Experimental results demonstrate that our method provides a more stringent measurement of group privacy risks, yielding a reliable assessment of the disparity in group privacy risks. Furthermore, to promote privacy protection fairness in DPML, we enhance the standard DP-SGD algorithm with an adaptive group-specific gradient clipping strategy, inspired by the design of canaries in differential privacy auditing studies. Extensive experiments confirm that our algorithm effectively reduces the disparity in group privacy risks, thereby enhancing the fairness of privacy protection in DPML. The code is avaliable at: https://github.com/ruayz/fair-privacy.
@@ 1 Introduction
Artificial intelligence (AI), particularly machine learning (ML), has been widely adopted across various sectors, augmenting and even replacing human decision-making. However, its growing integration in critical domains like healthcare, finance, and judiciary has raised critical concerns, including data privacy breaches, algorithmic biases, lack of explainability, security vulnerabilities etc. [huang2022overview]. Among these ethical issues and risks, privacy and fairness have emerged as two pivotal and widely discussed challenges [ekstrand2018privacy], attracting substantial attention from the research community.
@@
Research in privacy protection and fairness has made significant strides independently, and the intersection of these two critical issues has also gained considerable attention. Some studies aim to achieve both privacy protection and outcome fairness simultaneously in ML models [xu2019achieving, jagielski2019differentially, ding2020differentially, tran2021differentially, lowy2023stochastic]. Other works explore how privacy mechanisms affect the outcome fairness [bagdasaryan2019differential] and propose methods to mitigate the unfairness introduced by such mechanisms [xu2021removing, tran2021fairness, esipova2022disparate]. However, whether AI systems can provide equal privacy protections to different groups is also a noteworthy yet understudied problem at the intersection of fairness and privacy. As highlighted in [ekstrand2018privacy], this raises an essential yet still underexplored and insufficiently addressed question: Do AI systems offer fair or equitable privacy protections across groups? This question raises both ethical and practical concerns, as certain groups may face disproportionately higher privacy leakage risks, violating principles of fairness and equality.
@@
To address this issue, it is essential first to provide a rigorous answer to the question. Prior studies have empirically examined whether privacy leakage risks are evenly distributed across groups, and their auditing methods for group privacy risk rely on averaging the performance across data points within each group under membership inference attacks (MIAs) [chang2021privacy, yaghini2022disparate]. The MIAs employed in these studies are formulated based on the membership inference game (MIG) introduced in [yeom2018privacy], which captures the average behavior across all data points. However, such an average-case MIG may obscure the heterogeneity of individual privacy risks within groups and potentially underestimate the privacy leakage risks faced by certain groups. This, in turn, may lead to inaccurate and unreliable measurements of inter-group disparities in privacy risk. Consequently, the issue of privacy inequality may not be adequately uncovered.
@@
Measuring. Therefore, we aim to provide a tighter measurement of the privacy risk of data points and then analyze the disparities in privacy risk across different groups. To this end, we can leverage the leave-one-out attack (LOOA), which is formulated based on the worst-case MIG proposed by [ye2022enhanced] and is capable of estimating the worst-case privacy leakage risk for each data point. However, the computational cost of LOOA is prohibitively high, rendering its practical implementation nearly infeasible. To address this challenge, we propose an approximate version of the worst-case MIG to efficiently audit the approximate worst-case privacy risk of individual data points. Our experiments demonstrate that: 1) The attack simulating our proposed MIG can achieve comparable performance to LOOA as the number of attack rounds increases; 2) The individual privacy risks evaluated by our method can be reliable for assessing group privacy risk.
@@
Then we define a fairness metric to quantify the degree of privacy unfairness (i.e., the disparity of privacy risk across groups). Through experiments, we clearly demonstrate that our auditing method significantly outperforms previous auditing methods under the same conditions. Specifically, our method reveals greater group privacy risk and more effectively captures privacy inequality. Consistent with prior research [chang2021privacy, yaghini2022disparate], we find that existing ML algorithms exhibit significant unfairness in privacy risks across groups. While differentially private ML (DPML) algorithms can bind the magnitude of privacy risk disparities between groups, a certain degree of disparity still persists.
@@
Mitigating. Upon providing a more thorough answer to the above question, we seek to alleviate this issue. Inspired by the design of canaries in DP auditing studies [nasr2023tight, annamalai2024nearly, steinke2023privacy], we confirm that groups with larger gradient norms under training process—indicating greater contributions to model updates—are more prone to higher privacy leakage risks. Building on this insight, we enhance the existing DPML algorithm by adaptively setting group-specific gradient clipping norms. Extensive experimental results demonstrate that our algorithm effectively mitigates the disparity of group privacy risk, promoting the ethical and effective deployment of AI systems.
@@ 2.1 Differential privacy
Differential Privacy (DP), proposed by [dwork2006calibrating], is a privacy framework designed to address privacy leakage. It has become the predominant method for ensuring algorithmic privacy [ponomareva2023dp]. In the following, we introduce the approximate ($\epsilon,\delta$)-DP definition.
Definition 1 ($(\epsilon,\delta)$-Differential Privacy [dwork2006our]).
@@
DP-SGD [abadi2016deep] is a widely adopted algorithm in DPML [ponomareva2023dp]. It integrates DP concepts with stochastic gradient descent (SGD). This integration ensures model privacy by employing gradient clipping and noise addition within the SGD framework, adhering to the $(\epsilon,\delta)$-DP definition. The pseudocode of DP-SGD is shown in Algo. 2.
@@ 2.2 Black-box member inference attacks
The goal of membership inference attacks (MIAs) is to determine whether a specific data record is part of the training dataset. We focus on a black-box setting, where the adversary only has access to model outputs, reflecting a more realistic scenario where the training process is inaccessible [annamalai2024nearly]. In black-box MIAs, the adversary infers membership by analyzing the model’s output behavior, typically using the sample’s output loss as an inference score or decision basis, relying on the observation that models tend to show smaller losses for training samples [yeom2018privacy].
@@
MIGs conceptualize MIAs as inference games between a privacy auditor (i.e., the adversary) and a challenger. MIAs are typically carried out by simulating the MIGs through multiple rounds of random experiment. Various definitions of MIGs have been proposed, each designed to capture different aspects of privacy risk [ye2022enhanced].
Most MIAs follow the average-case MIG framework [yeom2018privacy] (see Def.6 in App.A)), which evaluates the vulnerability of a target model to the adversary, emphasizing the average behavior across data points [ye2022enhanced]. A common strategy formalized under this framework is the global attack (GA), where a single inference threshold is determined based on the aggregate behavior of all data points in a given round [yeom2018privacy, sablayrolles2019white]. The group-based attack (GBA) extends GA by assigning a distinct threshold to each group, enabling a more fine-grained analysis of group-level privacy risks and offering deeper insight into disparities across demographic partitions [chang2021privacy, yaghini2022disparate].
Nonetheless, such average-case MIAs fail to capture worst-case privacy risks for individual data records. The worst-case MIG  [ye2022enhanced] (see Def.7 in App.A) addresses this limitation by evaluating the maximum risk a single record may encounter. A concrete instance is the Leave-One-Out Attack (LOOA), which independently evaluates each data point’s worst-case exposure, aligning closely with the principles of DP. However, LOOA is computationally intensive, as it requires evaluating each record separately.
@@
Privacy auditing is designed to proactively assess privacy risks and quantify potential leakage, typically during the model development phase. In contrast, MIAs are conducted post-deployment by adversaries aiming to exploit trained models. Privacy auditing uses MIAs for evaluation but with more background knowledge for the adversary, including access to the original dataset and knowledge of the optimal threshold. This setup simulates worst-case scenarios, enabling rigorous assessment of privacy leakage risks [chang2021privacy]. In this work, we focus on privacy auditing to systematically evaluate privacy vulnerabilities. Privacy auditing using the GA (PA-GA)[yeom2018privacy], the GBA (PA-GBA) [chang2021privacy, yaghini2022disparate], and the LOOA (PA-LOOA) [nasr2023tight, annamalai2024nearly] for evaluating privacy leakage risks is detailed in Algos. 34, and 5 of App. A, respectively.
@@ 2.3 The fairness of privacy
Few studies have explored fairness in the context of privacy, and those that do vary in their research focus. For instance, one study finds that fairness-aware algorithms can exacerbate disparities in privacy leakage across groups [chang2021privacy]. Meanwhile, other research highlights that ML algorithms exhibit significant group-level disparities in privacy leakage, and that DPML algorithms can help reduce these disparities [yaghini2022disparate]. These auditing mechanisms typically rely on average-case attacks, which assess privacy risks based on the average behavior of data points. While effective at capturing general trends, such approaches may overlook the nuanced privacy risks faced by individual samples, potentially concealing disparities between groups.
In our work, we aim to more precisely evaluate group-level privacy risks compared to prior studies. To ensure fair comparison, we adopt the same attack metrics as in [chang2021privacy, yaghini2022disparate], including inference scores based on per-example loss and attack success rates.
@@ 3.1 Approximate worst-case privacy auditing
We validate our hypothesis through practical experiments using the widely used MNIST dataset, training a Convolutional Neural Network (CNN) with SGD. Due to computational limitations, we randomly select 60 samples per class, totaling 600 samples to audit for PA-LOOA and PA-ALOOA. We evaluate the attacker’s performance using the accuracy metric (i.e., attack success rate) from [shokri2017membership], which measures the agreement between the adversary’s guesses and the actual status. Instead of evaluating overall accuracy, we compute individual accuracy for each data point. Detailed experimental settings and additional results are provided in the App. B.
@@ 3.2 Definition of group privacy risk parity
We assess the privacy leakage risk of data points using the attacker’s membership advantage, following prior work [yaghini2022disparate, yeom2018privacy]. Below, we formally define the notion of individual privacy risk (IPR) as applied to a single data record in this study.
@@
This formulation quantifies the adversary’s normalized advantage over random guessing. Building on the concept of IPR, we can extend it to define Group Privacy Risk (GPR) as in  [chang2021privacy].
@@ 3.3 Comparison with privacy auditing by average-case attacks
We compare the GPR and the GPRP metrics measured by our auditing method, PA-ALOOA, with those obtained from privacy auditing by average-case attacks (PA-ACAs) in previous studies: PA-GA [yaghini2022disparate] and PA-GBA [chang2021privacy]. To ensure a fair comparison, all three methods are configured identically, keeping the training dataset and model consistent across each repeated experiment. Specifically, for PA-ALOOA, we set the number of audit samples to $m=n$, following the same setting used in PA-ACAs. This setup ensures consistency and reflects a real-world scenario in which the privacy risk of every training sample is audited. The key distinction lies in threshold determination: PA-ACAs computes thresholds based on the aggregate behavior of multiple data points, whereas PA-ALOOA assigns a unique threshold to each sample, based on its behavior across all repeated experiments.
@@
Consistent with prior studies on privacy and fairness [bagdasaryan2019differential, xu2021removing, esipova2022disparate], our foundational analysis focuses on the MNIST dataset. Detailed experimental configurations and supplementary results for other datasets are included in the App. C. The results of the GPR and GPRP metrics on the MNIST dataset are presented in Fig. 3 and Tab. 1, with all values reported in percentage points for clarity. As shown in Fig. 3, the three auditing methods exhibit consistent patterns in the distribution of privacy risks across groups. Among them, PA-ALOOA consistently yields significantly higher GPR values across all model architectures compared to the other two auditing methods. Specifically, Tab. 1 shows that for the CNN model, the GPRP value obtained by PA-GA and PA-GBA suggests that DP-SGD results in higher disparity than standard SGD. This observation contradicts the conclusion in [yaghini2022disparate], which asserts that DPML algorithms should be able to bind the disparity of GPR relative to non-private counterparts. Such inconsistency implies that PA-ACAs underestimate the GPR, leading to inaccurate measurements and incorrect conclusions. In summary, PA-ALOOA offers a more rigorous means of capturing privacy risk, and thus providing more reliable evaluations of privacy unfairness across groups.
@@ 4.1 Experimental observations
In DP auditing literature, canaries are often created as mislabeled samples [nasr2023tight, annamalai2024nearly, steinke2023privacy]. These samples generate larger gradient values during model training, which contribute more to parameter updates, thereby increasing the likelihood of being memorized by the model. Motivated by the design of canaries, we hypothesize that during training, the larger a group’s contribution to the gradient, the more likely the model is to memorize that group. Thus, groups with larger contributions are expected to face a higher privacy leakage risk compared to those with smaller contributions.
@@ 4.2 Design of mitigation algorithm for DP-SGD
As shown in Algo. 1, our proposed algorithm, DP-SGD-Scale (abbreviated as DP-SGD-S), differs from the standard DP-SGD in Lines 6–14. In each iteration, it estimates the relative contribution of each group’s samples to the overall gradient and uses this information to adaptively adjust the clipping bound for each group. To preserve privacy, we add noise to the clipped group-level gradient statistics used for computing the group-specific clipping bounds $C^{k}$. This additional privacy cost is incorporated into the overall privacy accounting via the composition theorem [abadi2016deep, xu2021removing]. Although different groups are assigned distinct clipping bounds, we conservatively bound the sensitivity of the final aggregated gradient by $\max_{k}C^{k}$, ensuring that the overall mechanism satisfies $(\epsilon,\delta)$-DP in the same sense as DP-SGD. Following prior work [xu2021removing, esipova2022disparate], we set $\sigma_{1}\approx 10\sigma_{2}$ so that the privacy cost of computing $C^{k}$ is negligible relative to the total privacy budget.